Sammy brought just enough fencing to border either rectangular plot or square plot the perimeters of the plots are the same how many meters of fencing did she buy?
(X + 2) meters on each side of the square
(3x + 2) meters horizontally on the rectangle
(X - 1) meters vertically on the rectangular

Answer by stanbon(75887)

Sammy brought just enough fencing to border either rectangular plot or square plot the perimeters of the plots are the same how many meters of fencing did she buy?
(X + 2) meters on each side of the square
Perimeter of square = 4(x+2) = 4x+8
-----------------------------
(3x + 2) meters horizontally on the rectangle
(X - 1) meters vertically on the rectangular
Perimeter of rectangle = 2(3x+2 + x-1) = 2(4x+1) = 8x+2
---------
Equation:
4x+8 = 8x+2
4x = 6
x = 3/2 meter
-----------------
Ans: 4x+8 = 4(3/2)+8 = 6+8 = 14 meters
-------------
Cheers,
Stan H.